What must be the distance between the point charge q 1 = 26.3C and q 2 = 38.2C in order to have a electrostatic attractive force between them has a magnitude of 4.89N.


1
Expert's answer
2021-01-29T20:01:29-0500

From the Coulomb's Law, we have:


"F=\dfrac{kQ_1Q_2}{r^2},""r=\sqrt{\dfrac{kQ_1Q_2}{F}},""r=\sqrt{\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot26.3\ C\cdot38.2\ C}{4.89\ N}}=1.36\cdot10^6\ m."